Eric is 1 mile ahead of Dave
Eric is driving 60 mph
Dave is driving 70 mph in the same direction
How many miles til Dave catches Eric?
Could you show me the formula for this question also?
This is not out of a book, a friend from work asked me this and told me the answer was 4.9999miles, and I wanted to make sure he was right. Thanks

Answer by edjones(8007)

d=st d=distance, s=speed, t=time
d+1=70t for Dave
d=60t for Eric.
.
60t+1=70t Replace 60t for d in the 1st equation.
10t=1
t=1/10 hr until Dave catches Eric.
In that .1 hr Dave goes 6 mi and Dave goes 7 mi
.
Ed